# Continuous-time q-learning in jump-diffusion models under Tsallis entropy

### v3

@@ 2.2 Weak convergence of sampled dynamics to relaxed version
$\mathbb{E}^{\mathbb{Q}}\left[I_{i}\right]=$ $\mathbb{E}^{\mathbb{Q}}\left[v(s_{i},X_{s_{i}})-v(s_{i-1},X_{s_{i-1}})\right]$
$=$ $\mathbb{E}^{\mathbb{Q}}\Bigg{[}\int_{s_{i-1}}^{s_{i}}\left(v_{t}(s,X_{s})+b(s,X_{s},u_{s_{i-1}})v_{x}(s,X_{s})+\frac{1}{2}\sigma^{2}(s,X_{s},u_{s_{i-1}})v_{xx}(s,X_{s})\right)ds$
$\quad+\int_{s_{i-1}}^{s_{i}}\int_{\mathds{R}}\left(v(s,X_{s}+\varphi(s,X_{s},u_{s_{i-1}},z))-v(s,X_{s})\right)\nu(dz)ds\Bigg{]}.$ (2.18)
As $\xi_{i-1}$ is independent of $X_{s_{i-1}}$, and $v$ satisfies the PDE (2.15), we have
$\mathbb{E}^{\mathbb{Q}}\Bigg{[}\int_{s_{i-1}}^{s_{i}}\left(v_{t}(s_{i-1},X_{s_{i-1}})+b(s_{i-1},X_{s_{i-1}},u_{s_{i-1}})v_{x}(s,X_{s})+\frac{1}{2}\sigma^{2}(s_{i-1},X_{s_{i-1}},u_{s_{i-1}})v_{xx}(s,X_{s})\right)ds$
$\quad+\int_{s_{i-1}}^{s_{i}}\int_{\mathds{R}}\left(v(s_{i-1},X_{s_{i-1}}+\varphi(s_{i-1},X_{s_{i-1}},u_{s_{i-1}},z))-v(s_{i-1},X_{s_{i-1}})\right)\nu(dz)ds\Bigg{]}$
$=\mathbb{E}^{\mathbb{Q}}\Bigg{[}\int_{s_{i-1}}^{s_{i}}\left(v_{t}(s_{i-1},X_{s_{i-1}})+\tilde{b}(s_{i-1},X_{s_{i-1}})v_{x}(s,X_{s})+\frac{1}{2}\tilde{\sigma}^{2}(s_{i-1},X_{s_{i-1}})v_{xx}(s,X_{s})\right)ds$
$\quad+\int_{s_{i-1}}^{s_{i}}\int_{U}\int_{\mathds{R}}\left(v(s_{i-1},X_{s_{i-1}}+\varphi(s_{i-1},X_{s_{i-1}},u,z))-v(s_{i-1},X_{s_{i-1}})\right)\nu(dz)\pi(u|s_{i-1},X_{s_{i-1}})ds\Bigg{]}$
@@
$\quad+\mathbb{E}^{\mathbb{Q}}\bigg{[}\int_{s_{i-1}}^{s_{i}}\int_{\mathds{R}}\left(v(s,X_{s}+\varphi(s,X_{s},u_{s_{i-1}},z))-v(s_{i-1},X_{s_{i-1}}+\varphi(s_{i-1},X_{s_{i-1}},u_{s_{i-1}},z))\right)\nu(dz)ds\bigg{]}$
$\quad-\mathbb{E}^{\mathbb{Q}}\bigg{[}\nu(\mathds{R})\int_{s_{i-1}}^{s_{i}}\left(v(s,X_{s})-v(s_{i-1},X_{s_{i-1}})\right)ds\bigg{]}$
@@
$:=\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(1)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(2)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(3)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(4)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(5)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(6)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(7)}\right].$ (2.20)
For $w=v,v_{t},v_{x},v_{xx},b(\cdot,\cdot,u_{s_{i-1}})$ or $\sigma^{2}(\cdot,\cdot,u_{s_{i-1}})$, we have from Itô’s rule that
@@
$\quad=\int_{s_{i-1}}^{\ell}\left(w_{t}(s,X_{s})+b(s,X_{s},u_{s_{i-1}})w_{x}(s,X_{s})+\frac{1}{2}\sigma^{2}(s,X_{s},u_{s_{i-1}})w_{xx}(s,X_{s})\right)ds$
$\qquad+\int_{s_{i-1}}^{s_{i}}\int_{\mathds{R}}\left(w(s,X_{s}+\varphi(s,X_{s},u_{s_{i-1}},z))-w(s,X_{s})\right)\nu(dz)ds$
$\qquad+\int_{s_{i-1}}^{\ell}\sigma(s,X_{s},u_{s_{i-1}})w_{x}(s,X_{s})dW_{s}$
$\qquad+\int_{s_{i-1}}^{s_{i}}\int_{\mathds{R}}\left(w(s,X_{s-}+\varphi(s,X_{s-},u_{s_{i-1}},z))-w(s,X_{s-})\right)\tilde{N}(ds,dz)$ (2.21)
with $\tilde{N}(ds,dz):=N(ds,dz)-\nu(dz)ds$. Using $w\in C^{2}_{k}$ and Assumption (A2), we deduce
$\left|\int_{\mathds{R}}\left(w(s,x+\varphi(s,x,u,z))-w(s,x)\right)\nu(dz)\right|\leq M_{k}\left|\int_{\mathds{R}}\left(|x+\varphi(s,x,u,z)|^{k}+|x|^{k}\right)\nu(dz)\right|$
$\qquad\leq M_{k}\left|\int_{\mathds{R}}\left(|\varphi(s,x,u,z)|^{k}+|x|^{k}\right)\nu(dz)\right|\leq M_{k}(1+|x|^{k}),$ (2.22)
where $M_{k}>0$ is a constant depending on $k$ that may be different from line to line.
Similar to the proof of Theorem 4.1 in Jia et al. (2025), it is not difficult to verify that $\mathbb{E}^{\mathbb{Q}}[|I_{i}^{(l)}|]$, $l=1,\ldots,6$, can be bounded by $C_{k}\|v\|_{C_{k}^{4}}(s_{i}-s_{i-1})^{2}$. For the term $I_{i}^{(7)}$, by using Itô’s rule to $v(\ell,Y_{\ell})$ with $Y_{\ell}=X_{\ell}+\varphi(\ell,X_{\ell},u_{s_{i-1}},z)$, it holds that
$v(\ell,X_{\ell}+\varphi(\ell,X_{\ell},u_{s_{i-1}},z))-v(s_{i-1},X_{s_{i-1}}+\varphi(s_{i-1},X_{s_{i-1}},u_{s_{i-1}},z))$
$\qquad=\int_{s_{i-1}}^{\ell}\left(v_{t}(s,Y_{s})+(1+\varphi_{x}(s,X_{s},u_{s_{i-1}},z))b(s,X_{s},u_{s_{i-1}})\right)ds$
$\qquad\quad+\int_{s_{i-1}}^{\ell}\frac{1}{2}(1+\varphi_{xx}(s,X_{s},u_{s_{i-1}},z))\sigma^{2}(s,X_{s},u_{s_{i-1}})ds$
$\qquad\quad+\int_{s_{i-1}}^{\ell}\int_{\mathds{R}}\left(v(s,Y_{s}+\varphi(s,Y_{s},u_{s_{i-1}},z)-v(s,Y_{s})\right)\nu(dz)ds$
$\qquad\quad+\int_{s_{i-1}}^{\ell}(1+\varphi_{x}(s,X_{s},u_{s_{i-1}},z))\sigma(s,X_{s},u_{s_{i-1}})dW_{s}$
$\qquad\quad+\int_{s_{i-1}}^{\ell}\int_{\mathds{R}}\left(v(s,Y_{s-}+\varphi(s,Y_{s-},u_{s_{i-1}},z)-v(s,Y_{s-})\right)\tilde{N}(ds,dz).$ (2.23)
By a similar discussion of $\mathbb{E}^{\mathbb{Q}}[|I_{i}^{(l)}|]$, $l=1,\ldots,6$ and (2.2), we have from Assumption (A4) and (2.2) that $\mathbb{E}^{\mathbb{Q}}[|I_{i}^{(7)}|]$ can be bounded by $C_{k}\|v\|_{C_{k}^{4}}(s_{i}-s_{i-1})^{2}$. Thus, $\mathbb{E}^{\mathbb{Q}}\left[I_{i}\right]\leq C_{k}\|v\|_{C_{k}^{4}}(s_{i}-s_{i-1})^{2}$. Then, the desired result follows from $\sum_{i=1}^{j}(s_{i}-s_{i-1})^{2}\leq T\|G_{0:T}\|$. ∎
@@
Let Assumptions (A1), (A2), (A4), (A5) and (A6) hold with $k\geq 2$. Given $(t,x)\in[0,T]\times\mathds{R}$ and $\pi\in\Pi_{t}$, consider the corresponding objective functionals $J^{0}$ given by (2.6) and J{J} given by (2.9). If $l_{p}(\pi(\cdot|\cdot,\cdot))\in C_{k}^{2,0}$, $\tilde{l}\in C_{k}^{2}$ and $\tilde{l}_{p}(t,\cdot)\in C_{k}^{4}(\mathds{R})$ for any $t\in[0,T]$, there exists a constant $C_{k}>0$ only depending on $(k,b,\sigma,\varphi,\nu,f,g,\tilde{b},\tilde{\sigma},\tilde{f},\tilde{l}_{p},\pi)$ such that $|J^{0}(t,x)-{J}(t,x)|\leq C_{k}\|G_{t:T}\|$ for all time grids $G_{t:T}$.
@@ 2.3 Exploratory HJB equation and policy improvement iteration
$V_{t}(t,x)+\sup_{\pi\in\mathcal{P}(U)}\Bigg{\{}V_{x}(t,x)\int_{U}b(t,x,u)\pi(u)du+\frac{1}{2}V_{xx}(t,x)\int_{U}\sigma^{2}(t,x,u)\pi(u)du$ (2.24)
$\quad+\int_{U}\int_{\mathds{R}}\left(V\left(t,x+\varphi(t,x,u,z)\right)-V(t,x)\right)\nu(dz)\pi(u)du+\int_{U}(f(t,x,u)+\gamma l_{p}(\pi(u)))\pi(u)du\Bigg{\}}=0$
with the terminal condition $V(T,x)=g(x)$ for all $x\in\mathds{R}$.
In order to characterize the optimal feedback policy, we introduce a scalar Lagrange multiplier $\psi(t,x):[0,T]\times\mathds{R}\to\mathds{R}$ to enforce the constraint $\int_{U}\pi(u)du=1$, and a Karush–Kuhn–Tucker (KKT) multiplier $\xi(t,x,u):[0,T]\times\mathds{R}\times U\to\mathds{R}_{+}$ to enforce the constraint $\pi(u)\geq 0$. The Lagrangian is written by
${\cal L}(t,x;\pi)$
$=\int_{U}\left(V_{x}(t,x)b(t,x,u)+\frac{\sigma^{2}(t,x,u)}{2}V_{xx}(t,x)+\int_{\mathds{R}}\left(V\left(t,x+\varphi(t,x,u,z)\right)-V(t,x)\right)\nu(dz)\right)\pi(u)du$
@@
    where, the Hamiltonian ${\cal H}(t,x,u,v)$ is defined as, for $(t,x,u)\in[0,T]\times\mathds{R}\times U$ and $v\in C^{1,2}([0,T)\times\mathds{R})\cap C([0,T]\times\mathds{R})$,
    ${\cal H}(t,x,u,v,v_{x},v_{xx})$ $:=b(t,x,u)v_{x}(t,x)+\frac{\sigma^{2}(t,x,u)}{2}v_{xx}(t,x)+f(t,x,u)$
    $\quad+\int_{\mathds{R}}\left(v\left(t,x+\varphi(t,x,u,z)\right)-v(t,x)\right)\nu(dz).$ (2.26)
    Then, it follows from the constraints on $\pi(u)\geq 0$ that
    $\xi(t,x,u)=(-{\cal H}(t,x,u,V,V_{x},V_{xx})-\psi(t,x))_{+},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ (x)_{+}:=\max\{x,0\}.$ (2.27)
@@
For the general case when $p>1$, the optimal policy characterized in (2.28) and (2.29) may no longer be a Gibbs measure, and particularly may not be Gaussian even in the linear quadratic control framework. In fact, the distribution of the optimal policy now heavily relies on the expression of the normalizing function $\psi(t,x)$ in (2.29). More importantly, the expression in (2.28) suggests that the density distribution of the optimal policy may not be supported on the whole $\mathds{R}$ in general, i.e., the sampled actions may concentrate on a compact set as the optimal policy is only defined on a compact subset of $\mathds{R}$; see the derived optimal policy distributions with compact support in Remarks 5.3 and Remark 5.6 in our two concrete examples.
The next result shows that the candidate optimal policy given by (2.28) and (2.30) leads to an improvement iteration rule. Recall the objective function $J(t,x;\pi)$ for a fixed admissible policy $\pi$ given by (2.9). Then, if the objective function $J(\cdot,\cdot;\pi)\in C^{1,2}([0,T)\times\mathds{R})\cap C([0,T]\times\mathds{R})$, it satisfies the following PDE:
$J_{t}(t,x;\pi)+J_{x}(t,x;\pi)\int_{U}b(t,x,u)\pi(u|t,x)du+\frac{1}{2}J_{xx}(t,x;\pi)\int_{U}\sigma^{2}(t,x,u)\pi(u|t,x)du$
$\qquad\qquad\qquad\qquad\qquad+\int_{U}\int_{\mathds{R}}\left(J\left(t,x+\varphi(t,x,u,z);\pi\right)-J(t,x;\pi)\right)\nu(dz)\pi(u|t,x)du$
@@
with the terminal condition $J(T,x:\pi)=g(x)$ for all $x\in\mathds{R}$.
@@
Let Assumptions (A1), (A2) and (A3) hold. For any given $\pi\in\Pi_{0}$, assume that the objective function $\tilde{J}(\cdot,\cdot;{\pi})\in C^{1,2}([0,T)\times\mathds{R})\cap C([0,T]\times\mathds{R})$ satisfies (2.3), and for $p>1$, there exists a function $\psi(t,x):[0,T]\times\mathds{R}\to\mathds{R}$ satisfying
@@
and ${\cal I}_{1}(\pi):=\lim_{p\downarrow 1}{\cal I}_{p}(\pi)=\frac{\exp\left\{\frac{1}{\gamma}{\cal H}(t,x,\cdot,J,J_{x},J_{xx})\right\}}{\int_{U}\exp\left\{\frac{1}{\gamma}{\cal H}(t,x,u,J,J_{x},J_{xx})\right\}du}$. Denote by $\pi^{\prime}={\cal I}_{p}(\pi)$ for $\pi\in\Pi_{0}$. If $\pi^{\prime}\in\Pi_{0}$, then $J\left(t,x;\pi^{\prime}\right)\geq J(t,x;\pi)$ for all $(t,x)\in[0,T]\times\mathds{R}$. Moreover, if the mapping ${\cal I}_{p}:\Pi_{0}\to\Pi_{0}$ has a fixed point $\pi^{*}\in\Pi_{0}$, then $\pi^{*}$ is the optimal policy that, for all $(t,x)\in[0,T]\times\mathds{R}$,
@@
$\sup_{\pi\in{\cal P}(U)}\int_{U}\left(q(u)\pi(u)-\frac{1}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du.$ (2.35)

### v4

@@ 2.2 Weak convergence of sampled dynamics to relaxed version
$\mathbb{E}^{\mathbb{Q}}\left[I_{i}\right]=$ $\mathbb{E}^{\mathbb{Q}}\Bigg[\int_{s_{i-1}}^{s_{i}}\left(v_{t}(s,X_{s})+b(s,X_{s},u_{s_{i-1}})v_{x}(s,X_{s})+\frac{1}{2}\sigma^{2}(s,X_{s},u_{s_{i-1}})v_{xx}(s,X_{s})\right)ds$
$\quad+\int_{s_{i-1}}^{s_{i}}\int_{\mathds{R}^{n}}\left(v(s,X_{s}+\varphi(s,X_{s},u_{s_{i-1}},z))-v(s,X_{s})\right)\lambda(s,X_{s},u_{s_{i-1}})\nu(dz)ds\Bigg].$ (2.19)
For simplicity, let us introduce the function $H^{v}(s,x,u):[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}$ given by
$H^{v}(s,x,u):=\int_{\mathds{R}^{n}}\left(v(s,x+\varphi(s,x,u,z))-v(s,x)\right)\lambda(s,x,u)\nu(dz).$
As $\xi_{i-1}$ is independent of $X_{s_{i-1}}$, and $v$ satisfies PDE (2.16), it holds that
$\mathbb{E}^{\mathbb{Q}}\Bigg[\int_{s_{i-1}}^{s_{i}}\left(v_{t}(s_{i-1},X_{s_{i-1}})+b(s_{i-1},X_{s_{i-1}},u_{s_{i-1}})v_{x}(s_{i-1},X_{s_{i-1}})\right)ds$
$\quad+\frac{1}{2}\int_{s_{i-1}}^{s_{i}}\sigma^{2}(s_{i-1},X_{s_{i-1}},u_{s_{i-1}})v_{xx}(s_{i-1},X_{s_{i-1}})ds+\int_{s_{i-1}}^{s_{i}}H^{v}(s_{i-1},X_{s_{i-1}},u_{s_{i-1}})ds\Bigg]$
$=\mathbb{E}^{\mathbb{Q}}\Bigg[\int_{s_{i-1}}^{s_{i}}\left(v_{t}(s_{i-1},X_{s_{i-1}})+\tilde{b}(s_{i-1},X_{s_{i-1}})v_{x}(s_{i-1},X_{s_{i-1}})\right)ds$
$\qquad+\frac{1}{2}\int_{s_{i-1}}^{s_{i}}\tilde{\sigma}^{2}(s_{i-1},X_{s_{i-1}})v_{xx}(s_{i-1},X_{s_{i-1}})ds+\int_{s_{i-1}}^{s_{i}}\int_{U}H^{v}(s_{i-1},X_{s_{i-1}},u)\pi(u|s_{i-1},X_{s_{i-1}})ds\Bigg]$
@@
$\quad+\mathbb{E}^{\mathbb{Q}}\bigg[\int_{s_{i-1}}^{s_{i}}\left(H^{v}(s,X_{s},u_{s_{i-1}})-H^{v}(s_{i-1},X_{s_{i-1}},u_{s_{i-1}})\right)ds\bigg]$
$:=\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(1)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(2)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(3)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(4)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(5)}\right]+\mathbb{E}^{\mathbb{Q}}\left[I_{i}^{(6)}\right].$ (2.21)
By using the fact $v\in C_{k}^{4}$ and Assumption (A4), it is straightforward to check that $H^{v}_{t},H^{v}_{x},H^{v}_{xx}$ have polynomial growth in $x$ and $u$. For $w=H,v_{t},v_{x},v_{xx},b(\cdot,\cdot,u_{s_{i-1}})$ or $\sigma^{2}(\cdot,\cdot,u_{s_{i-1}})$, we have from Itô’s rule that
@@
$=\int_{s_{i-1}}^{\ell}\left(w_{t}(s,X_{s})+b(s,X_{s},u_{s_{i-1}})w_{x}(s,X_{s})+\frac{1}{2}\sigma^{2}(s,X_{s},u_{s_{i-1}})w_{xx}(s,X_{s})\right)ds$
$\quad+\int_{s_{i-1}}^{\ell}\int_{\mathds{R}^{n}}\left(w(s,X_{s}+\varphi(s,X_{s},u_{s_{i-1}},z))-w(s,X_{s})\right)\lambda(s,X_{s},u_{s_{i-1}})\nu(dz)ds$
$\quad+\int_{s_{i-1}}^{\ell}\sigma(s,X_{s},u_{s_{i-1}})w_{x}(s,X_{s})dW_{s}$ (2.22)
$\quad+\int_{s_{i-1}}^{\ell}\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\left(w(s,X_{s-}+\varphi(s,X_{s-},u_{s_{i-1}},z))-w(s,X_{s-})\right){\bf 1}_{\{r\leq\lambda(s,X_{s-},u_{s_{i-1}})\}}\tilde{N}(ds,dz,dr)$
with $\tilde{N}(ds,dz,dr):=N(ds,dz,dr)-dr\nu(dz)ds$ being the compensated Poisson random measure. Using the fact $w\in C^{2}_{k}$ and Assumption (A2), we deduce that
$\left|\int_{\mathds{R}^{n}}\left(w(s,x+\varphi(s,x,u,z))-w(s,x)\right)\lambda(s,x,u)\nu(dz)\right|\leq M_{k}\left|\int_{\mathds{R}^{n}}\left(1+|x|^{k}\right)|\varphi(s,x,u,z)|\lambda(s,x,u)\nu(dz)\right|$
$\hskip 18.49988pt\leq M_{k}\left(1+|x|^{k}\right)\left|\int_{\mathds{R}^{n}}|\varphi(s,x,u,z)|\lambda(s,x,u)\nu(dz)\right|\leq M_{k}(1+|x|^{k}+|u|^{k}),$ (2.23)
where $M_{k}>0$ is a constant depending on $k$ that may vary from line to line.
Similar to the proof of Theorem 4.1 in Jia et al. (2025), by using (2.2) and (2.2), it is not difficult to verify that $\mathbb{E}^{\mathbb{Q}}[|I_{i}^{(l)}|]$ for $l=1,\ldots,6$ can be bounded by $C_{k}\|v\|_{C_{k}^{4}}(s_{i}-s_{i-1})^{2}$. Thus, we have $\mathbb{E}^{\mathbb{Q}}\left[I_{i}\right]\leq C_{k}\|v\|_{C_{k}^{4}}(s_{i}-s_{i-1})^{2}$. Then, the desired result follows from $\sum_{i=1}^{j}(s_{i}-s_{i-1})^{2}\leq T\|G_{0:T}\|$. ∎
@@
Let Assumptions (A1), (A2), (A4), (A5) and (A6) hold with $k\geq 2$. Given $(t,x)\in[0,T]\times\mathds{R}^{n}$ and $\pi\in\Pi_{t}$, consider the corresponding objective functionals $J^{0}$ given by (2.6) and J{J} given by (2.9). If $l_{p}(\pi(\cdot|\cdot,\cdot))\in C_{k}^{2,0}$, $\tilde{l}\in C_{k}^{2}$ and $\tilde{l}_{p}(t,\cdot)\in C_{k}^{4}(\mathds{R}^{n})$ for any $t\in[0,T]$, there exists a constant $C_{k}>0$ only depending on $(k,b,\sigma,\varphi,\nu,f,g,\tilde{b},\tilde{\sigma},\tilde{f},\tilde{l}_{p},\pi)$ such that $|J^{0}(t,x)-{J}(t,x)|\leq C_{k}\|G_{t:T}\|$ for all time grids $G_{t:T}$.
@@ 2.3 Exploratory HJB equation and policy improvement iteration
$V_{t}(t,x)+\sup_{\pi\in\mathcal{P}(U)}\Bigg\{$ $V_{x}(t,x)\cdot\int_{U}b(t,x,u)\pi(u)du+\frac{1}{2}\int_{U}\sigma^{2}(t,x,u)\pi(u)du\circ V_{xx}(t,x)$ (2.24)
$\quad+\int_{U}\int_{\mathds{R}^{n}}\left(V\left(t,x+\varphi(t,x,u,z)\right)-V(t,x)\right)\lambda(t,x,u)\nu(dz)\pi(u)du$
$\quad+\int_{U}(f(t,x,u)+\gamma l_{p}(\pi(u)))\pi(u)du\Bigg\}=0$
with the terminal condition $V(T,x)=g(x)$ for $x\in\mathds{R}^{n}$.
In order to characterize the optimal feedback policy, we introduce a scalar Lagrange multiplier $\psi(t,x):[0,T]\times\mathds{R}^{n}\to\mathds{R}$ to enforce the constraint $\int_{U}\pi(u)du=1$, and a Karush–Kuhn–Tucker (KKT) multiplier $\xi(t,x,u):[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}_{+}$ to enforce the constraint $\pi(u)\geq 0$. The Lagrangian is written by
${\cal L}(t,x;\pi)$ $=\int_{U}\left(V_{x}(t,x)\cdot b(t,x,u)+\frac{1}{2}\sigma^{2}(t,x,u)\circ V_{xx}(t,x)\right)\pi(u)du$
$\quad+\int_{U}\int_{\mathds{R}^{n}}\left(V\left(t,x+\varphi(t,x,u,z)\right)-V(t,x)\right)\lambda(t,x,u)\nu(dz)\pi(u)du$
@@
    where, the Hamiltonian ${\cal H}(t,x,u,v)$ is defined as, for $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$ and $v\in C^{1,2}([0,T)\times\mathds{R}^{n})\cap C([0,T]\times\mathds{R}^{n})$,
    ${\cal H}(t,x,u,v,v_{x},v_{xx})$ $:=b(t,x,u)\cdot v_{x}(t,x)+\frac{1}{2}\sigma^{2}(t,x,u)\circ v_{xx}(t,x)+f(t,x,u)$
    $\quad+\int_{\mathds{R}^{n}}\left(v\left(t,x+\varphi(t,x,u,z)\right)-v(t,x)\right)\lambda(t,x,u)\nu(dz).$ (2.26)
    Then, it follows from the constraints on $\pi_{p}^{*}(u)\geq 0$ that
    $\mathcal{H}(t,x,u,V,V_{x},V_{xx})+\psi(t,x)+\xi(t,x,u)\geq 0.$ (2.27)
    As $\xi(t,x,u)$ is the KKT multiplier for the non‑negativity constraint, it satisfies $\xi(t,x,u)\geq 0$ and the complementary slackness condition $\xi(t,x,u)\,\pi^{*}_{p}(u)=0$. If $\mathcal{H}(t,x,u,V,V_{x},V_{xx})+\psi(t,x)\geq 0$, we may set $\xi(t,x,u)=0$ and (2.27) holds; if $\mathcal{H}(t,x,u,V,V_{x},V_{xx})+\psi(t,x)<0$, then (2.27) forces $\xi(t,x,u)\geq-(\mathcal{H}(t,x,u,V,V_{x},V_{xx})+\psi(t,x))>0$. However, complementary slackness implies that whenever $\xi(t,x,u)>0$, we must have $\pi^{*}_{p}(u)=0$. Hence, in this case, we must have
    $\mathcal{H}(t,x,u,V,V_{x},V_{xx})+\psi(t,x)+\xi(t,x,u)=0.$
    Combining the two cases yields that
    $\xi(t,x,u)=(-{\cal H}(t,x,u,V,V_{x},V_{xx})-\psi(t,x))_{+},\penalty 10000\ \penalty 10000\ \text{with}\penalty 10000\ (x)_{+}:=\max\{x,0\}.$ (2.28)
@@
For the general case when $p>1$, the optimal policy characterized in (2.29) and (2.30) may no longer be a Gibbs measure, and particularly may not be Gaussian even in the linear quadratic control framework. In fact, the distribution of the optimal policy now heavily relies on the expression of the normalizing function $\psi(t,x)$ in (2.30). More importantly, the expression in (2.29) suggests that the density distribution of the optimal policy may not be supported on the whole $U$ in general, i.e., the sampled actions may concentrate on a compact set as the optimal policy is only defined on a compact subset of $U$; see the derived optimal policy distributions with compact support in Remarks 5.3 and Remark 5.6 in two concrete examples.
The next result shows that the candidate optimal policy given by (2.29) and (2.31) leads to an improvement iteration rule. Recall the objective function $J(t,x;\pi)$ for a fixed admissible policy $\pi$ given by (2.9). Then, if the objective function $J(\cdot,\cdot;\pi)\in C^{1,2}([0,T)\times\mathds{R}^{n})\cap C([0,T]\times\mathds{R}^{n})$, it satisfies the following PDE:
$J_{t}(t,x;\pi)+J_{x}(t,x;\pi)\cdot\int_{U}b(t,x,u)\pi(u|t,x)du+\frac{1}{2}\int_{U}\sigma^{2}(t,x,u)\pi(u|t,x)du\circ J_{xx}(t,x;\pi)$
$\qquad\qquad\qquad+\int_{U}\int_{\mathds{R}^{n}}\left(J\left(t,x+\varphi(t,x,u,z);\pi\right)-J(t,x;\pi)\right)\lambda(t,x,u)\nu(dz)\pi(u|t,x)du$
@@
with the terminal condition $J(T,x:\pi)=g(x)$ for all $x\in\mathds{R}^{n}$.
@@
Let Assumptions (A1), (A2) and (A3) hold. For any given $\pi\in\Pi_{0}$, assume that the objective function $\tilde{J}(\cdot,\cdot;{\pi})\in C^{1,2}([0,T)\times\mathds{R}^{n})\cap C([0,T]\times\mathds{R}^{n})$ satisfies (2.3), and for $p>1$, there exists a function $\psi(t,x):[0,T]\times\mathds{R}^{n}\to\mathds{R}$ satisfying
@@
and ${\cal I}_{1}(\pi):=\lim_{p\downarrow 1}{\cal I}_{p}(\pi)=\frac{\exp\left\{\frac{1}{\gamma}{\cal H}(t,x,\cdot,J,J_{x},J_{xx})\right\}}{\int_{U}\exp\left\{\frac{1}{\gamma}{\cal H}(t,x,u,J,J_{x},J_{xx})\right\}du}$. Denote by $\pi^{\prime}={\cal I}_{p}(\pi)$ for $\pi\in\Pi_{0}$. If $\pi^{\prime}\in\Pi_{0}$, then $J\left(t,x;\pi^{\prime}\right)\geq J(t,x;\pi)$ for all $(t,x)\in[0,T]\times\mathds{R}^{n}$. Moreover, if the mapping ${\cal I}_{p}$ has a fixed point $\pi^{*}\in\Pi_{0}$, then $\pi^{*}$ is the optimal policy that, for all $(t,x)\in[0,T]\times\mathds{R}^{n}$,
@@
$\sup_{\pi\in{\cal P}(U)}\int_{U}\left(q(u)\pi(u)+\frac{\gamma}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du.$ (2.36)